OpenStudy (anonymous):

The region between the graphs of x=y^2 and x=3y is rotated around the line y=3. Find the volume

OpenStudy (anonymous):

?

OpenStudy (jamesj):

Ok, so you need first to draw a diagram of the region. Where are the end-points?

OpenStudy (anonymous):

i cant tell

OpenStudy (anonymous):

it looks something like this 

OpenStudy (jamesj):

yes exactly so the origin and what's the other pt?

OpenStudy (jamesj):

It's when x = 3y and x = y^2. i.e., when 3y = y^2

OpenStudy (anonymous):

9?

OpenStudy (jamesj):

Hence x = 9; i.e., the other end point is (9,3)

OpenStudy (jamesj):

Now you're rotating about the line y = 3, so you need to measure distance from that line. You're going to want to the integral \[ 2\pi \int_0^9 (some \ function)^2 \ dx \]

OpenStudy (jamesj):

You need now to figure out that function.

OpenStudy (jamesj):

Here's one piece of it. The line x = 3y, or y = x/3, is for a given x a distance (3 - x/3) from the line y = 3. Now figure out the distance of the other curve from the line y = 3. The function you're squaring is ( (distance to curve #1) - (distance to curve #2) )

OpenStudy (anonymous):

are you sure its 2 pi?